OpenStudy (amyna):

Find limit of a sequence. Please help i don't know where to start! (1-(1/n))^(n)

OpenStudy (amyna):



OpenStudy (amyna):

and i haven't learned l'hopitals rule, so i can't use it in this problem

OpenStudy (freckles):

\[e^k=\lim_{n \rightarrow \infty}(1+\frac{k}{n})^n \text{ is a well known limit }\]

OpenStudy (amyna):

yes but how to i use it?

OpenStudy (freckles):

in place of k you have -1...

OpenStudy (amyna):

since its a minus instead of a plus

OpenStudy (freckles):

1-something is the same as 1+(-something)

OpenStudy (solomonzelman):